Vcc =\oV Re=8K.r No loon Rsig IE=ImA --VEE B=lbo Find re, vatage goin, max and min Vo

RL is also 8k
Answer #1

given IE=IMA, re- 17 - K= 26MN v des 266 DT=8+1) re = 101 X26= 2.63km hubrid-T model capacitor shorted Buvccd Vee gloended BVo=-Bib CROHRL) =-100 (8k/18k) ib 5-100(BX* K) ib 8t8 Vo. = -100 x 4000 i input loop KUL in DT b + iell00) + V = 0 le=íb + Bi